Suppose A and B are 3 × 3 matrices. Prove that
(i) (A + B)^T = B^T + A^T

(ii) det (AB)^T = det (A). det (B)
1
Expert's answer
2020-09-03T16:53:20-0400

Solution.

(i) Let A = (aij )n×n , AT = (a'ij )n×n , B = (bij )n×n , BT = (b'ij )n×n , (A+B) = (cij ) and (A+B) T = (dij ). Then dij = cji = aji +bji = (a'ij +b'ij ). Therefore, (A + B) T = BT +AT

(ii) det(AB)T = detAT. detBT

Hence, detAT= detA, detBT= detB

Therefore, det(AB)T= det(A). det(B)